OpenStudy (anonymous):

i just need someone to check the answer on this problem i got x=.3090 and x=-.8090 i used the quadratic formula and the problem is : 4x^2=1-2x

OpenStudy (amistre64):

4x^2 = 1 -2x 4x^2 +2x -1 = 0 \[\frac{-2}{2(4)} \pm\frac{\sqrt{4-4(-1)(4)}}{2(4)}\]

OpenStudy (amistre64):

\[{-1\over4}\pm \frac{\sqrt{20}}{8}\] \[-\frac{1}{4}\pm\frac{2\sqrt{5}}{8}\]

OpenStudy (amistre64):

\[\frac{-1\pm \sqrt{5}}{4}\] is what I get.....

OpenStudy (amistre64):

0.3090 .. is abt one of them so Id say you did good ;)

OpenStudy (amistre64):

-0.8090 is the other

OpenStudy (anonymous):

yay! lol thanx i wasnt sure!